OpenStudy (xbeenchantedx):

Graph each equation. State its domain and range. Scale all graphs by one. 1. f(x) = e^x 2. g(x) = ln x I need to find the coordinates to each individual equation as well as its domain and range. Please explain in as much detail as you can with how to solve for coordinates and find its domain and range.

OpenStudy (anonymous):

solve for what?

OpenStudy (anonymous):

one function is the inverse of the other that is for sure

OpenStudy (xbeenchantedx):

Solve for points of the function and the domain and range of the functions

OpenStudy (imtiaz7):

I think by taking limit approaches to infinity and we get to me is same that is e^x and for lnx =1

OpenStudy (anonymous):

so \[f\circ g(x)=e^{\ln(x)}=x\] and \[g\circ f(x)=\ln(e^x)=x\]

OpenStudy (anonymous):

what does "solve for points" on the function mean? i am not trying to be dense, but this makes no sense to me the domain of \(f(x)=e^x\) is \(\mathbb{R}\) and the range is \((0,\infty)\)

OpenStudy (anonymous):

since they are inverses, switch to get the domain and range of \(g(x)=\ln(x)\)

OpenStudy (anonymous):

Are they separate? You have to find range, domain of the functions: 1/ f(x) =e^x 2/ g(x) = ln x Am I right?

OpenStudy (xbeenchantedx):

First, I meant that you have to solve for coordinates on a graph of each question. Second, f(x) = e^x and g(x) = ln x are separate problems.

OpenStudy (xbeenchantedx):

I need to solve for each individual problem for its coordinates and domain and range as well as graphing it.

OpenStudy (xapproachesinfinity):

you mean sketching the graph? are you?

OpenStudy (anonymous):

oh you can't really solve for coordinates by hand, but the range and domain can be shown via limits.

OpenStudy (xbeenchantedx):

I need the coordinates to plot a set of points on the graph. Sorry for my vague question.

OpenStudy (anonymous):

except f(0) and g(1) which are 1 and 0 respectively

OpenStudy (xapproachesinfinity):

I think you mean plotting some point so that you can sketch the graph of each of e^x and lnx

OpenStudy (anonymous):

the only coordinate that people know offhand is e^0 and ln(1)

OpenStudy (xapproachesinfinity):

the problem is we don't know what is he exactly asking for?

OpenStudy (xbeenchantedx):

That's as much information that was given, which wasn't very useful to me or to anyone.

OpenStudy (xapproachesinfinity):

well just try to use calculator for some points problem is e is a rational number that just approximated so usualy to approch the graph of e^x in calculus or algebra we squeze it between two other function 2^x and 3^x so we get that graph of e^x which is somewhere between those two function that we know more about